AnimeGhoul8863:

Part A: Factor 5x2a2 − 19xa2 − 4a2. Show your work. (4 points) Part B: Factor x2 + 14x + 49. Show your work. (3 points) Part C: Factor x2 − 100. Show your work. (3 points)

AnimeGhoul8863:

Dis one is fairly easy so give me a sec and you can double check my work

Hero:

Part A: a^2 is common to each term Part B: Is a perfect square Part C: Is difference of squares

Hero:

We'll see what you do

AnimeGhoul8863:

Part A: factor out a^2 a^2 [5x^2 - 19x - 4] then we factor again a^2(5x+1)(x-4)

AnimeGhoul8863:

Part B: x2 + 14x + 49 (x^3+7x)+ (7x+49) =x(x+7) + 7(x+7) (x+7) (x+7) (x+7)^2

AnimeGhoul8863:

Part C: x2 − 100. x2 - 10^2 =(x+10)(x-10)

AnimeGhoul8863:

^^am i correct

AnimeGhoul8863:

@Hero

AnimeGhoul8863:

SOoooo am i e.e

AnimeGhoul8863:

@Hero xD THE HECK

Hero:

Looks right.

AnimeGhoul8863:

thx ill post the last question